40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 987     Entry time: Wed Sep 24 17:57:04 2008
Author: Alberto 
Type: Update 
Category: General 
Subject: ABSL: FSS Slow Actuator Control 
Rana, Alberto

Today when I started working with the PLL that I use to control the secondary laser on the ABSL experiment, I found that the beat between the two lasers was at a much higher temperature of NPRO than usual (about one Celsius Degrees higher, 49.79 instead of 48.7). It turned out that the main beam frequency had changed, and so had its frequency, because of a too much high value of the slow actuator gain on the FSS. We looked at the trend for the gain and noticed it had changed from 0.3 to 3 at about noon today. We brought it back to the old value and also optimized the single gains in the FSS slow servo to obtain a faster and stabler response to step changes in the laser temperature.

It is very important for the ABSL experiment that the frequency and the NPRO temperature of the main laser do not change.

** update:
you asked for:   diff 2008/09/25,0:00 2008/09/25,8:50:19 utc 'FSS[-_]SLOW'
LIGO controls: differences, 2008 09/25 00:00:00 utc vs. 2008 09/25 08:50:19 utc
__Epics_Channel_Name______   __Description__________   __value1____     __value2____
C1:PSL-FSS_SLOWKD                                      0.000000         0.001000
C1:PSL-FSS_SLOWKI                                     -0.001000        -0.001700
C1:PSL-FSS_SLOWKP                                     -0.000300        -0.001000

It seemed later that it was not being cool with the derivative gain up at -0.001, so I set it to zero. We really need some documentation on this
loop (e.g. pseudo code and a PID tuning procedure). Note that the PID record as documented in the EPICS Reference Manual
has been deprecated and so we run a perl script that Tobin wrote.
ELOG V3.1.3-